# Quantum Entanglement without nonlocal causation in (3,2)-dimensional spacetime

### v4

@@ II.2 Extending Bohm-Bub’s model to entangled particles
according to the initial sign of the terms in the first parenthesis of the r.h.s. of each one of the equations above, it is immediately evident that if the first particle collapses to the state |+ket|$\rangle$| + ⟩ the second one collapses to |ket|$\rangle$| - ⟩ and viceversa. In analogy with the original Bohm-Bub’s model, the $\xi^{(j)}_{\pm}(t)$ are assumed to be hidden random variables, almost everywhere non vanishing, and constant during measurement. The non-locality is here expressed by the fact that even if the two particles are separated by a space-like interval, the values of the quantities $R^{(j)}$ have an instantaneous mutual influence on the respective wavefunction collapses. Hereafter we introduce the following working hypothesis which is the theoretical core that is being considered. We assume that the correlated wavefunctions-collapses of the two particles - that now have their own individuality - are driven by an exchange of information mediated by a hidden sub-quantum field ${\mathscr{X}}({\bf x},t,\tau)$ which propagates through an extra temporal dimension described by the variable $\tau$. This hypothesis necessarily requires the extra time dimension because a superluminal exchange of information is forbidden in our standard space-time. We are replacing the hidden variables $\xi^{(j)}_{\pm}(t)$ with a completely different physical entity, that is, a real hidden physical field ${\mathscr{X}}({\bf x},t,\tau)$, almost nowhere vanishing (that is with the possible exception of a zero measure subset of space), a condition that could be ensured for example by the existence of a stochastic background component of the field. Let us remark that the variables $\xi_{i}(t)$ of the original Bohm-Bub model, as well as the field ${\mathscr{X}}({\bf x},t,\tau)$ introduced here, play a role only during the nonlinear, non-unitary dynamics of the state reduction, thus they do not belong to the ”family” of hidden variables theories where quantum mechanics is assumed to be a statistical approximation of some unknown deterministic theory where the values of the observables are already defined and fixed by unknown, hidden variables.
We tentatively assume that the dynamical evolution of the scalar field ${\mathscr{X}}({\bf x},t,\tau)$ is described by a wave equation
@@
accounting for information propagation at speed $w$ by means of ${\mathscr{X}}({\bf x},t,\tau)$ through the dimension $\tau$, and where the inhomogeneous source term $F[\psi_{\pm}^{(j)}(t),{\mathscr{X}}({\bf x},t,\tau)]$ should be chosen so that equation (19) fulfils the following requirements: i) it describes the propagation of the field ${\mathscr{X}}({\bf x},t,\tau)$ in the form of a non-dispersive impulse carrying the information of the outcome of a measure performed at some spatial location; ii) this information is isotropically propagated in space; iii) the information impulse carried by the field ${\mathscr{X}}({\bf x},t,\tau)$ is not attenuated with distance, and this assumption is required to comply with standard quantum non-locality which is independent of the distance between entangled entities. So we update equation (19) as
$\nabla^{2}{\mathscr{X}}(r,t,\tau)-\frac{1}{w^{2}}\frac{\partial^{2}{\mathscr{X}}(r,t,\tau)}{\partial\tau^{2}}\ +\ \sin[{\mathscr{X}}(r,t,\tau)]+{\widetilde{F}}[{\bf x},\psi_{\pm}^{(j)}(t),\tau;t_{0},\tau_{0}]=0$ (20)
@@
${\widetilde{R}}^{(1)}_{\pm}=\frac{|\psi^{(1)}_{\pm}(t)|^{2}}{|{\mathscr{X}}({\bf x}_{2}-{\bf w}\Delta\tau,t_{0},\tau_{0}+\Delta\tau)|^{2}}\ ,\quad\quad\quad{\widetilde{R}}^{(2)}_{\pm}=\frac{|\psi^{(2)}_{\pm}(t)|^{2}}{|{\mathscr{X}}({\bf x}_{1}-{\bf w}\Delta\tau,t_{0},\tau_{0}+\Delta\tau)|^{2}}$ (27)
@@
$R^{(1)}_{\pm}=\frac{|\psi^{(1)}_{\pm}(t)|^{2}}{|{\mathscr{X}}({\bf x}_{1},t_{0},\tau_{0}+\Delta\tau)|^{2}}\ ,\quad\quad\quad R^{(2)}_{\pm}=\frac{|\psi^{(2)}_{\pm}(t)|^{2}}{|{\mathscr{X}}({\bf x}_{2},t_{0},\tau_{0}+\Delta\tau)|^{2}}$ (28)
@@
Notice that the field ${\mathscr{X}}({\bf x},t,\tau)$ is assumed to propagate at finite (unknown) velocity $w$ according to equation (20) through the extra time $\tau$ only, and it is assumed to propagate at infinite velocity in our standard (3+1) spacetime. In fact, the denominators $|{\mathscr{X}}({\bf x}_{1,2}-{\bf w}\Delta\tau,t_{0},\tau_{0}+\Delta\tau)|^{2}$ entering the above defined functions ${\widetilde{R}}^{(1)}_{\pm}$ and ${\widetilde{R}}^{(2)}_{\pm}$ account for the propagation of ${\mathscr{X}}({\bf x},t,\tau)$ only in the extra time dimension because the usual time enters solely with $t_{0}$, the instant of measurement, thus implicitly assuming nonlocal instantaneous correlation in $t$ with the absence of a finite propagation speed. This is an important point to ensure the no-signalling condition discussed in the Introduction.
@@
The final outcome of the measure is determined by the evolution of $|\Phi^{(1)}\rangle|\Phi^{(2)}\rangle$ which depends on the values that the field ${\mathscr{X}}({\bf x},t_{0},\tau_{0}+\Delta\tau)$ had immediately before the measurement at ${\bf x}_{1}$ and ${\bf x}_{2}$, and thus it depends on the random component of the field and on the unknown details about the shaping of the information-carrying impulses. All this makes the individual outcomes of measurements on entangled entities unpredictable, outcomes which however remain perfectly correlated. In other words, although it is demonstrated that the Bell inequality is violated by non-product states gisin , nevertheless, also with the assumption done in the present work that a Bell state factorizes when both entangled entities interact simultaneously with the respective measurement devices, the physical information about the entanglement - and thus the violation of Bell inequality - is maintained by the collapse equations entailing correlated outcomes of the measures via the field ${\mathscr{X}}({\bf x},t,\tau)$.
@@ II.3 Proposal for a nonconventional experiment
$R^{(A,E)}_{\pm}=\frac{|\psi^{(A,E)}_{\pm}(t)|^{2}}{|{\mathscr{X}}({\bf x}_{A,E},t_{0},\tau_{0}+\Delta_{2}\tau)|^{2}}$ (34)
${\widetilde{R}}^{(A)}_{\pm}=\frac{|\psi^{(A)}_{\pm}(t)|^{2}}{|{\mathscr{X}}({\bf x}_{E}-\textbf{w}\Delta_{2}\tau,t_{0},\tau_{0}+\Delta_{2}\tau)|^{2}}\ ,\quad\quad\quad{\widetilde{R}}^{(E)}_{\pm}=\frac{|\psi^{(E)}_{\pm}(t)|^{2}}{|{\mathscr{X}}({\bf x}_{A}-\textbf{w}\Delta_{2}\tau,t_{0},\tau_{0}+\Delta_{2}\tau)|^{2}}$ (35)

### v5

@@ II.2 Extending Bohm-Bub’s model to entangled particles
according to the initial sign of the terms in the first parenthesis of the r.h.s. of each one of the equations above, it is immediately evident that if the first particle collapses to the state |+ket|$\rangle$| + ⟩ the second one collapses to |ket|$\rangle$| - ⟩ and viceversa. In analogy with the original Bohm-Bub’s model, the $\xi^{(j)}_{\pm}(t)$ are assumed to be hidden random variables, almost everywhere non vanishing, and constant during measurement. The non-locality is here expressed by the fact that even if the two particles are separated by a space-like interval, the values of the quantities $R^{(j)}$ have an instantaneous mutual influence on the respective wavefunction collapses. Hereafter we introduce the following working hypothesis which is the theoretical core that is being considered. We assume that the correlated wavefunctions-collapses of the two particles - that now have their own individuality - are driven by an exchange of information mediated by a hidden sub-quantum field ${\mathscr{X}}({\bf x},t,\tau)$ which propagates through an extra temporal dimension described by the variable $\tau$. This hypothesis necessarily requires the extra time dimension because a superluminal exchange of information is forbidden in our standard space-time. We are replacing the hidden variables $\xi^{(j)}_{\pm}(t)$ with a completely different physical entity, that is, a real hidden physical vector field ${\mathscr{X}}({\bf x},t,\tau)$, of components ${\mathscr{X}}_{\pm}({\bf x},t,\tau)$, almost nowhere vanishing (that is with the possible exception of a zero measure subset of space), a condition that could be ensured for example by the existence of a stochastic background component of the field. Let us remark that the variables $\xi_{i}(t)$ of the original Bohm-Bub model, as well as the field ${\mathscr{X}}({\bf x},t,\tau)$ introduced here, play a role only during the nonlinear, non-unitary dynamics of the state reduction, thus they do not belong to the ”family” of hidden variables theories where quantum mechanics is assumed to be a statistical approximation of some unknown deterministic theory where the values of the observables are already defined and fixed by unknown, hidden variables.
We tentatively assume that the dynamical evolution of the field ${\mathscr{X}}({\bf x},t,\tau)$ is described by a wave equation
@@
accounting for information propagation at speed $w$ by means of ${\mathscr{X}}({\bf x},t,\tau)$ through the dimension $\tau$, and where the inhomogeneous source term $F[\psi_{\pm}^{(j)}(t),{\mathscr{X}}({\bf x},t,\tau)]$ should be chosen so that equation (19) fulfils the following requirements: i) it describes the propagation of the field ${\mathscr{X}}({\bf x},t,\tau)$ in the form of a non-dispersive impulse carrying the information of the outcome of a measure performed at some spatial location; ii) this information is isotropically propagated in space; iii) the information impulse carried by the fields ${\mathscr{X}}_{\pm}({\bf x},t,\tau)$ is not attenuated with distance, and this assumption is required to comply with standard quantum non-locality which is independent of the distance between entangled entities. So we update equation (19) as
$\nabla^{2}{\mathscr{X}}_{\pm}(r,t,\tau)-\frac{1}{w^{2}}\frac{\partial^{2}{\mathscr{X}}_{\pm}(r,t,\tau)}{\partial\tau^{2}}\ +\ \sin[{\mathscr{X}}_{\pm}(r,t,\tau)]+{\widetilde{F}}[{\bf x},\psi_{\pm}^{(j)}(t),\tau;t_{0},\tau_{0}]=0$ (20)
@@
${\widetilde{R}}^{(1)}_{\pm}=\frac{|\psi^{(1)}_{\pm}(t)|^{2}}{|{\mathscr{X}}_{\pm}({\bf x}_{2}-{\bf w}\Delta\tau,t_{0},\tau_{0}+\Delta\tau)|^{2}}\ ,\quad\quad\quad{\widetilde{R}}^{(2)}_{\pm}=\frac{|\psi^{(2)}_{\pm}(t)|^{2}}{|{\mathscr{X}}_{\pm}({\bf x}_{1}-{\bf w}\Delta\tau,t_{0},\tau_{0}+\Delta\tau)|^{2}}$ (27)
@@
$R^{(1)}_{\pm}=\frac{|\psi^{(1)}_{\pm}(t)|^{2}}{|{\mathscr{X}}_{\pm}({\bf x}_{1},t_{0},\tau_{0}+\Delta\tau)|^{2}}\ ,\quad\quad\quad R^{(2)}_{\pm}=\frac{|\psi^{(2)}_{\pm}(t)|^{2}}{|{\mathscr{X}}_{\pm}({\bf x}_{2},t_{0},\tau_{0}+\Delta\tau)|^{2}}$ (28)
@@
Notice that the field ${\mathscr{X}}({\bf x},t,\tau)$ is assumed to propagate at finite (unknown) velocity $w$ according to equation (20) through the extra time $\tau$ only, and it is assumed to propagate at infinite velocity in our standard (3+1) spacetime. In fact, the denominators $|{\mathscr{X}}_{\pm}({\bf x}_{1,2}-{\bf w}\Delta\tau,t_{0},\tau_{0}+\Delta\tau)|^{2}$ entering the above defined functions ${\widetilde{R}}^{(1)}_{\pm}$ and ${\widetilde{R}}^{(2)}_{\pm}$ account for the propagation of ${\mathscr{X}}({\bf x},t,\tau)$ only in the extra time dimension because the usual time enters solely with $t_{0}$, the instant of measurement, thus implicitly assuming nonlocal instantaneous correlation in $t$ with the absence of a finite propagation speed. This is an important point to ensure the no-signalling condition discussed in the Introduction.
@@
The final outcome of the measure is determined by the evolution of $|\Phi^{(1)}\rangle|\Phi^{(2)}\rangle$ which depends on the values that the fields ${\mathscr{X}}_{\pm}({\bf x},t_{0},\tau_{0}+\Delta\tau)$ had immediately before the measurement at ${\bf x}_{1}$ and ${\bf x}_{2}$, and thus it depends on the random component of the field and on the unknown details about the shaping of the information-carrying impulses. All this makes the individual outcomes of measurements on entangled entities unpredictable, outcomes which however remain perfectly correlated. In other words, although it is demonstrated that the Bell inequality is violated by non-product states gisin , nevertheless, also with the assumption done in the present work that a Bell state factorizes when both entangled entities interact simultaneously with the respective measurement devices, the physical information about the entanglement - and thus the violation of Bell inequality - is maintained by the collapse equations entailing correlated outcomes of the measures via the field ${\mathscr{X}}({\bf x},t,\tau)$.
@@ II.3 Proposal for a nonconventional experiment
$R^{(A,E)}_{\pm}=\frac{|\psi^{(A,E)}_{\pm}(t)|^{2}}{|{\mathscr{X}}_{\pm}({\bf x}_{A,E},t_{0},\tau_{0}+\Delta_{2}\tau)|^{2}}$ (34)
${\widetilde{R}}^{(A)}_{\pm}=\frac{|\psi^{(A)}_{\pm}(t)|^{2}}{|{\mathscr{X}}_{\pm}({\bf x}_{E}-\textbf{w}\Delta_{2}\tau,t_{0},\tau_{0}+\Delta_{2}\tau)|^{2}}\ ,\quad\quad\quad{\widetilde{R}}^{(E)}_{\pm}=\frac{|\psi^{(E)}_{\pm}(t)|^{2}}{|{\mathscr{X}}_{\pm}({\bf x}_{A}-\textbf{w}\Delta_{2}\tau,t_{0},\tau_{0}+\Delta_{2}\tau)|^{2}}$ (35)